Write the following as an algebraic expression then simplify
a) The sum of three consecutive integers if x is the first consectutive integers.
b)The perimeter of a square with side length 3x +1.

Answer by ankor@dixie-net.com(12704)

Write the following as an algebraic expression then simplify
:
a) The sum of three consecutive integers if x is the first consecutive integers.
Sum = x + (x+1) + (x+2)
Sum = 3x + 3
Sum = 3(x+1)
:
b)The perimeter of a square with side length 3x +1.
P = 4(3x + 1)